OpenStudy (anonymous):

Find the area of the triangle with the given measurements. Round the solution to the nearest hundredth if necessary. B = 104°, a = 11 cm, c = 18 cm

OpenStudy (anonymous):

@lochana @bradely

OpenStudy (bradely):

A =(1/2)acsinB Plug into the values to get area Source: http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

96.06 cm^2?

OpenStudy (bradely):

A = (1/2)(11)(18)sin104 =99sin104 =96.06 cm^2

OpenStudy (anonymous):

thanks!